Let A = {(x, y, z) E R^3ㅣx - y + z = 0} where E represents "an element of" and R the set of real numbers
A mapping f is defined in R^(2*2) , the set of all square matrices of order 2 with entries in R by f: R^(2*2) → R^2
A → Av where v = (-1 2).
(a) Show that f is a linear mapping.
1
Expert's answer
2021-02-12T10:36:11-0500

From the question,

We'll show that f is linear if for "A_1, A_2 \in R^{2*2}, \alpha, \beta \in R," we have that "f(\alpha A_1+\beta A_2)= \alpha f(A_1)+\beta f(A_2)" To this end,

Let "A_1= \begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}, A_2=\begin{pmatrix}
 e & f \\
 g & h
\end{pmatrix}\\
So, \\
\alpha A_1+ \beta A_2= \alpha \begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}+ \beta \begin{pmatrix}
 e & f \\
 g & h
\end{pmatrix}\\
\implies\\
\alpha A_1+ \beta A_2=\begin{pmatrix}
 \alpha a+ \beta e & \alpha b+ \beta f\\
 \alpha c +\beta g & \alpha a+ \beta h
\end{pmatrix}\\
Now, \\
f(\alpha A_1+ \beta A_2)=(\alpha A_1+ \beta A_2)v\\
f(\alpha A_1+ \beta A_2)=
\begin{pmatrix}
 \alpha a+ \beta e & \alpha b+ \beta f\\
 \alpha c +\beta g & \alpha a+ \beta h
\end{pmatrix}\begin{pmatrix}
 -1 \\
 2
\end{pmatrix}\\
f(\alpha A_1+ \beta A_2)=\begin{pmatrix}
 -\alpha a- \beta e +2 \alpha b+2 \beta f\\
 -\alpha c -\beta g +2 \alpha a+2 \beta h
\end{pmatrix}\\
f(\alpha A_1+ \beta A_2)=\begin{pmatrix}
 -\alpha a +2 \alpha b \\
 \alpha c +2 \alpha d
\end{pmatrix}+\begin{pmatrix}
 -\beta e +2 \beta f \\
 -\beta g +2 \beta h
\end{pmatrix}\\
f(\alpha A_1+ \beta A_2)= \alpha\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}\begin{pmatrix}
 -1 \\
 2
\end{pmatrix}+\beta \begin{pmatrix}
 e & f \\
 g & h
\end{pmatrix}\begin{pmatrix}
 -1 \\
 2
\end{pmatrix}\\
f(\alpha A_1+ \beta A_2)= \alpha A_1v+ \beta A_2v\\
f(\alpha A_1+ \beta A_2)=\alpha f(A_1)+\beta f(A_2)\\

\text{Thus, f is a linear mapping}"